OpenStudy (anonymous):

What is the value of 10C7? i really dont get this

OpenStudy (jkristia):

looks like a hex number.

OpenStudy (phi):

do you mean 10 choose 7 , often written as \[\left(\begin{matrix}10 \\ 7\end{matrix}\right)\] if so, this is the same as \[ \frac{10!}{(10-7)!7!} \] more commonly written as \[ \frac{10!}{3! 7!} \]